OpenStudy (mitu12):

A 6ft tall tent standing next to a cardboard box casts a 9 foot shadow. If the cardboard box casts a shadow that is 6ft long then how tall is it?

OpenStudy (tkhunny):

It's a proportion: \(\dfrac{Tent}{Tent\;Shadow} = \dfrac{Box}{Box\;Shadow}\)

OpenStudy (mitu12):

x=4

OpenStudy (tkhunny):

Are you guessing?

OpenStudy (mitu12):

thats the answer i got

OpenStudy (tkhunny):

Is it right?

OpenStudy (mitu12):

i think so

OpenStudy (anonymous):

How can the cardboard box cast a shadow of both 9 and 6 feet?

OpenStudy (tkhunny):

I was hoping for a better response. If you formalize your notations and produce work that can be followed and repeated, you can know for sure. If x is the height of the box, then \(\dfrac{6}{9} = \dfrac{x}{6}\). Solving for x, we have x = (6*6)/9 = 36/9 = 4

OpenStudy (mitu12):

thanks

OpenStudy (tkhunny):

English Class. When you remove a prepositional phrase, there should be a complete sentence left. A 6ft tall tent standing next to a cardboard box casts a 9 foot shadow. Identify Prepositional Phrase [] A 6ft tall tent [standing next to a cardboard box] casts a 9 foot shadow. Remove Prepositional Phrase A 6ft tall tent casts a 9 foot shadow.